OpenStudy (anonymous):

Two mechanics worked on a car. The first mechanic worked for 20 hours, and the second mechanic worked for 15 hours. Together they charged a total of $3000 . What was the rate charged per hour by each mechanic if the sum of the two rates was $175 per hour?

OpenStudy (snowsurf):

This best way to solve this is write two equations. Lets mechanic that work for 20 hours be x and the other for 15 be y. X and y are their rates in dollar per hour so we have an equation of 20x + 15y = 3000 And we know their rates total to 175, so we have x + y = 175. So your two equations are 20x + 15y = 3000 x + y = 175 Now you can use any method whether substitution or elimination to solve for x and y.

OpenStudy (anonymous):

How would you use the substitution in this method?

OpenStudy (snowsurf):

You would solve either for x or y and then substitute into the other equation.